2. You read the classified ads in your newspaper and found a company for sale. The asking price was $400,000. However, you further read that the company was in legal trouble with the government and would require you to spend approximately $250,000 (legal fee) after tax to bring the company on l (or back to life). The company would be expected to generate after tax cash flows of $100-009 a year. 1oo With a required rate ofreturn of 12%, is it a good buy if (Please explain your answer with your calculations)
Answer #1

Cost to buy the company = asking price+ legal fees

=4,00,000+2,50,000

=6,50,000$

It is expected to generate 150000 per year

Hence present value of cash inflow = 150000/12%

=1250,000$

Thus net benefit = 1250,000-650000 = 600,000$

Thus he should purchase the company